OpenStudy (anonymous):

Find the radius of a sphere whose volume is 288 Π cm3. Round to the nearest tenth.

OpenStudy (anonymous):

cm^3**

OpenStudy (jfraser):

What's the equation for the volume of a sphere? Rearrange the equation to solve for r instead of V

OpenStudy (anonymous):

Find the radius of a sphere whose volume is 288 Π cm^3. Round to the nearest tenth.

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ 288\pi}}\ cm^3\qquad \textit{volume of a sphere}={\color{brown}{ \cfrac{4}{3}\pi r^3}}\qquad thus \\ \quad \\ 288\pi = \cfrac{4}{3}\pi r^3\qquad \textit{solve for }r\)

OpenStudy (anonymous):

Okay, give me a minute to solve.

OpenStudy (anonymous):

6 cm.

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

I need help on a different question.

OpenStudy (anonymous):

@jdoe0001 Find the volume of a sphere with a surface area of 324 π in.2

OpenStudy (jdoe0001):

\(\bf 324\pi in^2\qquad \textit{surface area of a sphere}=4\pi r^2 \\ \quad \\ 324\pi = 4\pi r^2\qquad \textit{solve for }r\qquad \textit{then use it }\to \cfrac{4}{3}\pi r^3 \iff volume\)

OpenStudy (anonymous):

For some reason, I didn't get the right answer.

OpenStudy (jdoe0001):

well... what did you get for "r"?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Do I put 9 into this: 9(pi)4/3(pi)r^3?

OpenStudy (jdoe0001):

hmm yeap... is 9... so \(\bf \cfrac{4}{3}\pi r^3\implies \cfrac{4}{3}\pi 9^3\implies 972\pi\approx 3053.6\)

OpenStudy (anonymous):

OH okay, I just didn't replace the r for 9.

OpenStudy (jdoe0001):

yes, because the same sphere with that surface area and thus radius, will also have the same radius appled to the volume

OpenStudy (anonymous):

I'm going to post another question so I could give you more medals for your help.

OpenStudy (jdoe0001):

tis easier if I dunno someone else may anyhow

OpenStudy (anonymous):

If a sphere has a radius 15 m, what is the volume of the hemisphere?